# Evaluating the Evolution of YOLO (You Only Look Once) Models: A Comprehensive Benchmark Study of YOLO11 and Its Predecessors

### v1

@@ IV-A2 Africa Wildlife Dataset
YOLOv9t and YOLOv9s consistently excel across all metrics, delivering high accuracy while maintaining small model sizes, low GFLOPs, and short inference times, as shown in Table VII, and Figures 13, 14, and 15. This demonstrates the robustness of YOLOv9’s smaller models and their effectiveness on small datasets. In contrast, YOLO5ux and YOLO11x show suboptimal accuracy despite their larger sizes and longer inference times, likely due to overfitting. Most large models underperformed on this dataset, with the exception of YOLOv10x, which benefited from a modern architecture that prevents overfitting.
IV-A3 Ships and Vessels Dataset
Table VIII presents the performance of YOLO models on the Ships and Vessels dataset, a large dataset featuring tiny objects with varying rotations. Overall, the models demonstrated moderate effectiveness in detecting ships and vessels, with mAP50-95 ranging from 0.273 to 0.327. This performance suggests that YOLO algorithms may face challenges in accurately detecting smaller objects, and the dataset’s diversity in object sizes and rotations provides a comprehensive test of the models’ capabilities in these conditions.
Accuracy:
The disparity between mAP50-95 and mAP50, illustrated in Figure 16, underscores the challenges YOLO models face with higher IoU thresholds when detecting small objects. Additionally, YOLO models struggle with detecting objects of varying rotations. Among the models, YOLO11x achieved the highest accuracy, with a mAP50 of 0.529 and a mAP50-95 of 0.327, closely followed by YOLO11l, YOLO11m, and YOLO11s, which recorded mAP50 values of 0.529, 0.528, and 0.53, and mAP50-95 values of 0.327, 0.325, and 0.325, respectively. These results highlight the robustness of the YOLO11 family in detecting small and tiny objects. In contrast, YOLOv3u-tiny, YOLOv8n, YOLOv3u, and YOLOv5n exhibited the lowest accuracy, with mAP50 scores of 0.489, 0.515, 0.519, and 0.514, and mAP50-95 scores of 0.273, 0.297, 0.298, and 0.298, respectively. This suggests the outdated architecture of YOLOv3u and the potential underfitting of smaller models due to the large dataset size.
Precision and Recall:
Figure 17 indicates that YOLOv5ux outperformed other models, achieving a precision of 0.668 and a recall of 0.555. It was closely followed by YOLOv9m (precision of 0.668, recall of 0.551) and YOLOv8m (precision of 0.669, recall of 0.525), both of which are significantly smaller in size (40.98 Mb for YOLOv9m and 52.12 Mb for YOLOv8m). In contrast, YOLO11n and YOLOv10s exhibited lower performance, with precisions of 0.574 and 0.586 and recalls of 0.51 and 0.511, respectively, likely due to underfitting issues. Generally, YOLO11 models tended to produce false positives, reflected in their low precision and high recall. Meanwhile, YOLOv10 underperformed in both precision and recall, despite being one of the newest models in the YOLO family.
Computational Efficiency:
As illustrated in Figures 18 and 19, YOLOv3u-tiny achieved the fastest processing time at 2 ms, closely followed by YOLOv8n and YOLOv5un, both recording 2.3 ms. YOLOv10 and YOLO11 models also excelled in speed, with YOLOv10n and YOLO11n achieving rapid inference times of 2.4 ms and 2.5 ms, along with GFLOPs counts of 8.2 and 6.3, respectively. In contrast, YOLOv9e exhibited the slowest speed, with an inference time of 7.6 ms and a GFLOPs count of 189.3, highlighting the trade-off between accuracy and efficiency within the YOLOv9 family.
Overall Performance:
The results in Table VIII and Figures 16, 17, and 18 demonstrate that YOLO11s and YOLOv10s excelled in accuracy while maintaining compact sizes, low GFLOPs, and quick processing times. In contrast, YOLOv3u, YOLOv8x, and YOLOv8l fell short of expectations despite their larger sizes and longer processing times. These findings highlight the robustness and reliability of the YOLO11 family, particularly in improving the YOLO family’s performance in detecting small and tiny objects while ensuring efficient processing. Additionally, the results reveal the underperformance of YOLOv9 models when faced with large datasets and small objects, despite their modern architecture.
@@ IV-A3 Ships and Vessels Dataset
| YOLO11x | 0.596 | 0.538 | 0.529 | 0.327 | 0.8 | 6.1 | 0.3 | 7.2 | 109 | 194.4 |
@@ IV-B Discussion
| Version |
| YOLOv3u-tiny | 28 | 1 | 6 | 11 | ||||||||
| YOLOv3u | 20 | 24 | 28 | 28 | ||||||||
| YOLOv5un | 27 | 6 | 2 | 4 | ||||||||
| YOLOv5us | 24 | 7 | 8 | 9 | ||||||||
| YOLOv5um | 17 | 15 | 13 | 18 | ||||||||
| YOLOv5ul | 14 | 19 | 21 | 23 | ||||||||
| YOLOv5ux | 17 | 27 | 26 | 27 | ||||||||
| YOLOv8n | 26 | 5 | 4 | 5 | ||||||||
| YOLOv8s | 23 | 9 | 11 | 10 | ||||||||
| YOLOv8m | 15 | 17 | 16 | 20 | ||||||||
| YOLOv8l | 13 | 22 | 22 | 22 | ||||||||
| YOLOv8x | 8 | 26 | 27 | 26 | ||||||||
| YOLOv9t | 20 | 12 | 3 | 1 | ||||||||
| YOLOv9s | 7 | 15 | 10 | 6 | ||||||||
| YOLOv9m | 4 | 21 | 15 | 15 | ||||||||
| YOLOv9c | 9 | 25 | 19 | 19 | ||||||||
| YOLOv9e | 12 | 28 | 24 | 25 | ||||||||
| YOLOv10n | 25 | 2 | 5 | 3 | ||||||||
| YOLOv10s | 19 | 3 | 9 | 7 | ||||||||
| YOLOv10m | 5 | 10 | 12 | 12 | ||||||||
| YOLOv10b | 9 | 12 | 18 | 14 | ||||||||
| YOLOv10l | 11 | 17 | 20 | 17 | ||||||||
| YOLOv10x | 2 | 22 | 23 | 21 | ||||||||
| YOLO11n | 22 | 3 | 1 | 2 | ||||||||
| YOLO11s | 16 | 8 | 7 | 8 | ||||||||
| YOLO11m | 1 | 11 | 14 | 13 | ||||||||
| YOLO11l | 3 | 14 | 17 | 16 | ||||||||
| YOLO11x | 5 | 19 | 25 | 24 |
The analysis of Table IX yields several critical observations:
@@ IV-B1 Accuracy
YOLO11m consistently emerged as a top performer, frequently ranking among the highest, closely followed by YOLOv10x, YOLO11l, YOLOv9m, and YOLO11x. This underscores the robust performance of the YOLO11 family across varying IoU thresholds and object sizes, which can be attributed to their use of C2PSA for the preservation of contextual information, leading to improved convergence and overall performance. In addition, the implementation of large-kernel convolutions and partial self-attention modules helped increase the performance of the algorithm.
@@
Conversely, YOLOv3u-tiny exhibited the lowest accuracy, particularly in the Africa Wildlife and Ships and Vessels datasets, with YOLOv5un and YOLOv8n showing slightly better but still sub-par results. This suggests that YOLO11 models are currently the most reliable for applications demanding high accuracy.
Closely following the performance of the YOLO11 family, the YOLOv9 models demonstrate their effectiveness in detecting objects across various sizes and different IoU thresholds. However, they may struggle with small objects, as seen in the Ships and Vessels dataset. In contrast, the YOLOv10 family, despite its later introduction, exhibited relatively lower accuracy in the Traffic Signs and Africa Animals datasets, resulting in an average accuracy drop of 2.075% compared to the YOLOv9 models in those datasets. The slight underperformance of YOLOv10 can be attributed to its adoption of the One-to-One Head approach instead of Non-Maximum Suppression (NMS) for defining bounding boxes. This strategy can struggle to capture objects effectively, particularly when dealing with overlapping items, as it relies on a single prediction per object. This limitation helps explain the relatively subpar results observed in the second dataset.
Similarly, the outdated architecture of YOLOv3u contributed to its inferior performance, averaging 6.5% lower accuracy than the YOLO11 models. This decline can be traced back to its reliance on the older Darknet-53 framework, first introduced in 2018, which may not adequately address contemporary detection challenges.
@@ IV-B2 Computational Efficiency
YOLOv10n consistently outperformed other models in terms of speed and GFLOPs count, ranking among the top across all three datasets in terms of speed and 5th in terms of GFLOPs count. YOLOv3u-tiny, YOLOv10s, and YOLO11n also demonstrated notable computational efficiency.
YOLOv9e exhibited the slowest inference times and a very high GFLOPs count across the datasets, illustrating the trade-off between accuracy and efficiency. YOLO11’s speed improvements, attributable to their use of the C3k2 block, make it suitable for applications where rapid processing is essential, surpassing YOLOv10 and YOLOv9 models, in terms of speed by %1.41 and %31 on average, respectively.
While YOLOv9 models excelled in accuracy, their inference times were among the slowest, making them less ideal for time-sensitive applications. In contrast, YOLOv10 models, though slightly slower than the YOLO11 variants, still offer a commendable balance between efficiency and speed. Their performance is well-suited for time-sensitive scenarios, providing rapid processing without significantly sacrificing accuracy, making them a viable option for real-time applications.
@@ IV-B3 Model Size
YOLOv9t was the smallest model, ranking first across all three datasets, followed by YOLO11n and YOLOv10n. This efficiency in model size underscores the advancements in newer YOLO versions, especially YOLOv10, showcasing the effectiveness of implementing the Spatial-Channel Decoupled Downsampling for efficient parameter utilization.
YOLOv3u was the largest model, highlighting its inefficiency compared to its more modern counterparts due to its outdated architecture.

### v2

@@ IV-A2 Africa Wildlife Dataset
YOLOv9t and YOLOv9s consistently excel across all metrics, delivering high accuracy while maintaining small model sizes, low GFLOPs, and short inference times, as shown in Table VII, and Figures 15 and 16. This demonstrates the robustness of YOLOv9’s smaller models and their effectiveness on small datasets. In contrast, YOLOv12x and YOLO5ux show suboptimal accuracy despite their larger sizes and longer inference times, likely due to overfitting. Most large models underperformed on this dataset, with the exception of YOLOv10x, which benefited from a modern architecture that prevents overfitting.
@@
| YOLO11x | 0.596 | 0.538 | 0.529 | 0.327 | 0.8 | 6.1 | 0.3 | 7.2 | 109 | 194.4 | ||||||||
| YOLOv12n | 0.571 | 0.505 | 0.505 | 0.307 | 1.1 | 4.7 | 0.4 | 6.2 | 5.34 | 6.3 | ||||||||
| YOLOv12s | 0.58 | 0.529 | 0.516 | 0.319 | 1.6 | 5 | 0.4 | 7 | 18.1 | 21.2 | ||||||||
| YOLOv12m | 0.584 | 0.543 | 0.525 | 0.325 | 1.1 | 4.6 | 0.4 | 6.1 | 39 | 67.1 | ||||||||
| YOLOv12l | 0.588 | 0.532 | 0.522 | 0.323 | 2.6 | 5.2 | 0.4 | 8.2 | 51.2 | 88.5 | ||||||||
| YOLOv12x | 0.586 | 0.538 | 0.52 | 0.321 | 1.2 | 9.3 | 0.4 | 10.9 | 114 | 198.5 |
IV-A3 Ships and Vessels Dataset
Table VIII presents the performance of YOLO models on the Ships and Vessels dataset, a large dataset featuring tiny objects with varying orientations. Overall, the models demonstrated moderate effectiveness in detecting ships and vessels, with mAP50-95 ranging from 0.273 to 0.327. This performance suggests that YOLO algorithms may face challenges in accurately detecting smaller objects with different orientations.
Accuracy:
The disparity between mAP50-95 and mAP50, illustrated in Figure 17, underscores the challenges YOLO models face with higher IoU thresholds when detecting small objects. Additionally, YOLO models struggle with detecting objects of varying rotations. Among the models, YOLO11x achieved the highest accuracy, with a mAP50 of 0.529 and a mAP50-95 of 0.327, closely followed by YOLO11l, YOLO11m, and YOLO11s, which recorded mAP50 values of 0.529, 0.528, and 0.53, and mAP50-95 values of 0.327, 0.325, and 0.325, respectively. These results highlight the robustness of the YOLO11 family in detecting small and tiny objects. In contrast, YOLOv3u-tiny, YOLOv8n, YOLOv3u, and YOLOv5n exhibited the lowest accuracy, with mAP50 scores of 0.489, 0.515, 0.519, and 0.514, and mAP50-95 scores of 0.273, 0.297, 0.298, and 0.298, respectively. This suggests the outdated YOLOv3u architecture and the potential underfitting of smaller models with large size of datasets.
Precision and Recall:
Figure 18 indicates that YOLOv5ux outperformed other models, achieving a precision of 0.668 and a recall of 0.555. It was closely followed by YOLOv9m (precision of 0.668, recall of 0.551) and YOLOv8m (precision of 0.669, recall of 0.525), both of which are significantly smaller in size (40.98 Mb for YOLOv9m and 52.12 Mb for YOLOv8m). In contrast, YOLOv12n and YOLO11n exhibited lower performance, with precisions of 0.571 and 0.574 and recalls of 0.505 and 0.51, respectively, likely due to underfitting issues. Generally, YOLOv12 and YOLO11 models tended to produce false positives, reflected in their low precision and high recall. Meanwhile, YOLOv10 underperformed in both precision and recall, despite being one of the newest models in the YOLO family.
Computational Efficiency:
As illustrated in Figure 19, YOLOv3u-tiny achieved the fastest processing time at 2 ms, closely followed by YOLOv8n and YOLOv5un, both recording 2.3 ms. YOLOv10 and YOLO11 models also excelled in speed, with YOLOv10n and YOLO11n achieving rapid inference times of 2.4 ms and 2.5 ms, along with GFLOPs counts of 8.2 and 6.3, respectively. In contrast, YOLOv12x exhibited the lowest speed, with 10.9 ms inference time and 198.5 GFLOPs count, highlighting the shortcomings of the newest YOLO family.
Overall Performance:
The results in Table VIII and Figures 17, 18, and 19 demonstrate that YOLO11s and YOLOv10s excelled in accuracy while maintaining compact sizes, low GFLOPs, and quick processing times. In contrast, YOLOv3u, YOLOv8x, and YOLOv8l fell short of expectations despite their larger sizes and longer processing times. These findings highlight the robustness and reliability of the YOLO11 family, particularly in improving the YOLO family’s performance in detecting small and tiny objects while ensuring efficient processing. Additionally, the results reveal the underperformance of YOLOv9 models when faced with large datasets and small objects, despite their modern architecture.
@@ IV-B Discussion
| Version | Acc. | Speed | GFLOPs | Size |
| YOLOv3u tiny | 33 | 1 | 7 | 33 |
| YOLOv3u | 25 | 28 | 33 | 13 |
| YOLOv5un | 32 | 6 | 3 | 5 |
| YOLOv5us | 29 | 8 | 10 | 11 |
| YOLOv5um | 21 | 15 | 15 | 21 |
| YOLOv5ul | 18 | 21 | 25 | 27 |
| YOLOv5ux | 22 | 31 | 31 | 32 |
| YOLOv8n | 31 | 5 | 5 | 6 |
| YOLOv8s | 26 | 9 | 13 | 12 |
| YOLOv8m | 19 | 18 | 19 | 24 |
| YOLOv8l | 17 | 23 | 26 | 26 |
| YOLOv8x | 11 | 30 | 32 | 31 |
| YOLOv9t | 23 | 14 | 4 | 1 |
| YOLOv9s | 10 | 16 | 12 | 7 |
| YOLOv9m | 5 | 22 | 18 | 18 |
| YOLOv9c | 13 | 26 | 23 | 23 |
| YOLOv9e | 16 | 33 | 28 | 30 |
| YOLOv10n | 30 | 2 | 6 | 4 |
| YOLOv10s | 23 | 3 | 11 | 8 |
| YOLOv10m | 9 | 9 | 14 | 14 |
| YOLOv10b | 11 | 12 | 22 | 17 |
| YOLOv10l | 14 | 16 | 24 | 20 |
| YOLOv10x | 2 | 26 | 27 | 25 |
| YOLO11n | 27 | 3 | 1 | 3 |
| YOLO11s | 19 | 6 | 8 | 10 |
| YOLO11m | 1 | 11 | 17 | 15 |
| YOLO11l | 2 | 13 | 20 | 19 |
| YOLO11x | 8 | 23 | 29 | 28 |
| YOLOv12n | 27 | 19 | 1 | 2 |
| YOLOv12s | 7 | 25 | 8 | 9 |
| YOLOv12m | 5 | 20 | 16 | 16 |
| YOLOv12l | 4 | 28 | 21 | 22 |
| YOLOv12x | 15 | 32 | 30 | 29 |
The analysis of Table IX yields critical observations:
@@ IV-B1 Accuracy
YOLO11m consistently emerged as a top performer, frequently ranking among the highest, closely followed by YOLOv10x, YOLO11l, YOLOv9m, and YOLO11x. This underscores the robust performance of the YOLO11 family across varying IoU thresholds and object sizes, which can be attributed to their use of C2PSA for the preservation of contextual information, leading to improved convergence and overall performance. The inclusion of two small kernel convolutions and partial self-attention modules in these models helped capture more spatial features, leading to better accuracy in detecting both large and small objects. Additionally, the fine-tuning of anchor boxes and multi-scale training has enhanced the detection of objects at different resolutions, which is crucial in achieving high accuracy across diverse datasets.
@@
Conversely, YOLOv3u-tiny exhibited the lowest accuracy, particularly in the Africa Wildlife and Ships and Vessels datasets. YOLOv5un and YOLOv8n show slightly better but still sub-par results. This suggests that YOLO11 models are currently the most reliable for applications demanding high accuracy.
The YOLOv3u-tiny model’s underperformance can be attributed to its older architecture, which lacks the optimizations found in newer YOLO models. YOLOv3u is based on the Darknet-53 framework, which, although efficient for its time, struggles to capture complex patterns and fine details required for high-accuracy detection. Its relatively shallow architecture and absence of more recent enhancements such as attention mechanisms or advanced feature fusion techniques make it less effective in handling intricate object detection tasks, especially in datasets with a wide variety of object sizes, like those in the Africa Wildlife and Ships and Vessels datasets.
In contrast, the YOLOv9 models show effectiveness in detecting objects across various sizes and IoU thresholds, but they still struggle with small objects, particularly in the Ships and Vessels dataset. The YOLOv9 family benefits from improved training strategies, such as anchor box refinement and better scale handling, but may still fall short in detecting small objects in crowded or cluttered scenes due to its reliance on larger-scale feature maps and a less efficient processing pipeline for fine-grained features.
The YOLOv10 family, despite its later introduction, exhibited relatively lower accuracy in the Traffic Signs and Africa Animals datasets, resulting in an average accuracy drop of 2.075% compared to the YOLOv9 models in those datasets. YOLOv10’s adoption of the One-to-One Head approach instead of Non-Maximum Suppression (NMS) for defining bounding boxes contributed to this decrease. The One-to-One Head approach relies on predicting a single bounding box per object, which can be a limitation in scenarios with overlapping or closely packed objects. NMS, on the other hand, is more effective at filtering out redundant predictions, which can improve detection accuracy in these types of datasets. Therefore, the One-to-One Head’s inability to handle multiple overlapping predictions adequately explains YOLOv10’s underperformance, especially in datasets like Traffic Signs, where multiple signs are often close together.
Although the YOLOv12 family includes strong models like YOLOv12l, which ranked 4th in accuracy, it falls short compared to previous iterations, particularly considering its status as the newest YOLO version. It shows a -2.72% decline compared to YOLOv9 in the African Animals dataset and a -1% decrease in accuracy compared to YOLO11 in the Ships and Vessels dataset. This decline in accuracy is largely due to architectural choices. R-ELAN and area attention, while innovative, compromised global dependency modeling, impacting accuracy. Additionally, the adjusted MLP ratio prioritized certain computational aspects over feature representation depth, limiting its ability to capture intricate patterns. Consequently, YOLOv12 underperformed compared to the more balanced YOLOv11 and YOLOv9 families, demonstrating that newer architectural changes did not yield the expected gains in accuracy.
Similarly, YOLOv8n, though it performed better than YOLOv3u, still showed lower accuracy than the top-performing models in specific datasets. This model demonstrated a good balance between speed and accuracy but struggled with intricate object recognition tasks. The simplicity of its architecture, without more advanced modules like self-attention or multi-scale training, likely held it back from achieving the high accuracy observed in models like YOLO11 and YOLOv9.
Lastly, the YOLO11 models stand out for their ability to consistently achieve the highest accuracy. The use of advanced mechanisms like C2PSA and two kernel convolutions allowed for improved feature extraction across multiple scales, enhancing their ability to detect objects of varying sizes and complexities. These models’ robust performance across a variety of datasets, including challenging ones like Africa Wildlife and Ships and Vessels, demonstrates their ability to handle a wide range of object detection tasks with exceptional accuracy.
@@ IV-B2 Computational Efficiency
YOLOv10n consistently outperformed other models in terms of speed and GFLOPs count, ranking among the top for speed across all datasets and 5th in terms of GFLOPs. Its use of CSPNet as backbone design and optimization strategies allowed for high processing speed without significantly compromising performance. YOLOv3u-tiny, YOLOv10s, and YOLO11n also demonstrated excellent computational efficiency, with their lightweight architectures enabling faster inference times and lower resource usage.
Despite its accuracy advantages, YOLOv12 ranked as the worst family in terms of processing time and second in GFLOPs count. This is largely due to its attention-centric architecture, which enhances accuracy but significantly increases computational complexity. The integration of Area Attention and R-ELAN, while effective in capturing spatial features, introduces substantial overhead in feature extraction and aggregation, leading to slower inference times. Additionally, the reliance on FlashAttention contributes to inefficient memory access patterns, further impacting processing speed. Although these architectural choices improve contextual feature extraction, they hinder computational efficiency, making YOLOv12 less suitable for real-time applications compared to its predecessors. This performance trade-off illustrates that while attention mechanisms can elevate accuracy, they come at the cost of increased latency and computational demands.
YOLOv5un, YOLOv5us, and YOLOv8n exhibited strong efficiency, providing a good balance between speed and accuracy. These models performed well in both inference times and GFLOPs, with YOLOv5un showing slightly better speed than YOLOv8n potentially due to the implementation of C2k (CSP with larger kernel sizes) in YOLOv8 architecture that increases accuracy but with a trade-off in computational efficiency. YOLOv8n, despite its good speed performance, had a larger size and computational cost compared to some other lightweight models.
While YOLOv3u-tiny ranked highly in speed, its performance in accuracy was subpar, particularly in datasets like Africa Wildlife and Ships and Vessels. This highlights a critical trade-off: while speed is crucial for real-time applications, it becomes irrelevant if it severely compromises accuracy. YOLOv3u-tiny’s lightweight architecture, while optimized for faster inference, failed to detect objects accurately across varied and complex datasets, making it unsuitable for applications where precision is key. On the other hand, YOLOv3u struggled in both speed and accuracy, ranking poorly in both categories. Its reliance on the outdated Darknet-53 backbone contributed to this underperformance, as it was unable to meet modern detection challenges. The model’s low ranking in both computational efficiency and accuracy makes it less viable compared to newer architectures, such as YOLOv11 and YOLOv10, which offer better performance across both metrics.
In contrast, YOLO11 models demonstrated significant speed improvements, thanks to the C3k2 block, which reduces computational overhead while maintaining high accuracy. These models were faster than YOLOv10 by 1.41% and YOLOv9 by 31% on average, making them ideal for real-time applications requiring rapid processing without sacrificing accuracy. While YOLOv9 models excelled in accuracy, their slower inference times made them less ideal for time-sensitive tasks. YOLOv10, slightly slower than YOLO11, offered a good balance between speed and accuracy, making it well-suited for real-time applications.
Lastly, YOLOv5 and YOLOv8 families also performed well, with YOLOv5un ranking near the top for speed and efficiency, making it a solid choice for applications where both quick processing and solid accuracy are required.
@@ IV-B3 Model Size
YOLOv9t emerged as the smallest model, ranking first across all three datasets in terms of model size thanks to its use of GELAN architecture leading to more efficient use of parameters. It is followed closely by YOLOv12n, YOLO11n, and YOLOv10n. Specifically, YOLOv10n benefits from the Spatial-Channel Decoupled Downsampling technique, which effectively reduces the model size while maintaining high accuracy, making it a highly efficient choice for resource-constrained environments. This technique allows for better utilization of parameters, enabling smaller models to perform similarly to larger counterparts.
In contrast, YOLOv3u was the largest model, ranking poorly in terms of both size and accuracy. This inefficiency stems from its reliance on the outdated Darknet-53 architecture, which is larger and more computationally expensive compared to more modern designs. The size of YOLOv3u further exacerbates its subpar performance, as the model requires more resources and longer inference times, without corresponding improvements in detection accuracy. Its larger size makes it less suitable for deployment in environments where computational resources are limited or where real-time processing is essential.
This comparison highlights the trend in newer YOLO versions, such as YOLOv9, YOLO10, and YOLO11, where optimizations have led to smaller, more efficient models without significant sacrifices in accuracy or speed. These advancements make them more attractive for modern applications, where both model size and computational efficiency are critical. However, despite being the latest addition, YOLOv12 did not follow this trend, ranking third in terms of model size and computational efficiency.